OpenStudy (anonymous):

what is the derivative of ln(sqrt(x^2)) ? ? Thanks in advance!

OpenStudy (m):

1/x?

OpenStudy (m):

\[\ln \sqrt{x ^{2}}\]

OpenStudy (m):

is that what you mean?

OpenStudy (anonymous):

ln(sqrt(x^2)) = ln x (d/dx) ln x = 1/x right?

OpenStudy (m):

yes

OpenStudy (anonymous):

\[(d / dx) \ln \sqrt{x^{2}}\]

OpenStudy (anonymous):

so its ok to go ahead and cancel the exponent with the square root and then just do ln (x) ?